Do not understand need answers
Consider the reaction Write the equilibrium o nt for this reaction in terms of the eqr 1rium constants, K, and K, for reactions a and b below a.) P(s) + 5/2 Cl2(g) PCIs(g) K b.) PCl (g) + C(g)PCls(g) Kb
media%2Fa8f%2Fa8fcee21-b378-41fb-94b5-02
Answer #1

First, note that there is no PCl5

so we need to get rid of PCl5

the best will be to invert (b) since it has PCl3 and we need it in th eproduct side

so

P + 5/2Cl2 = PCl5 Ka

PCl5 = PCl3 + Cl2 1/Kb

add both equaitons

PCl5 + P + 5/2Cl2 = PCl5+  PCl3 + Cl2 Ka * 1/Kb

cancel common terms

P + 5/2Cl2 = PCl3 + Cl2 K total = Ka * 1/Kb

Q2.

now...

Krequired = [H2]^0.5 * [Br]^0.5 / [HBr]

from the data:

for

i)

H2 + Br = 2HBr

this is: inverted, and multiplied by 4

so

Keffect = (1/K)^4

ii)

this is: just inverted so

Keffect = 1/K

iii)

this is:

multiplied just by 2 so

Keffect = K^2